OpenStudy (anonymous):

is this correct (6, 3) and (8, 6). Slope (m) = change in y over change in x.? m= y2-y1/x2-x1

OpenStudy (anonymous):

6 - 8 = -2 3 - 6 = -3 -2/-3 = m (Slope)